OpenStudy (jenessa_2413):

A proportional relationship between the number of pounds of potatoes (x) and the price in dollars (y) is graphed, and the ordered pair (4, 3) is on the graphed line. Part A: What is the price of 1 pound of potatoes? Show your work, including the proportion you used to determine the price. (8 points) Part B: What does the ordered pair (8, 6) on the graph represent? Explain in words.

OpenStudy (jenessa_2413):

@welshfella

OpenStudy (jenessa_2413):

@Jaynator495 @whpalmer4

OpenStudy (jenessa_2413):

@robtobey2

OpenStudy (jenessa_2413):

@ganeshie8

OpenStudy (jenessa_2413):

@ParthKohli

OpenStudy (carrionbeast):

From what I know a proportion is a ratio(fraction) in this case we can model the proportion as for every four pounds of potatoes it costs three dollars. The opposite can be said too. For every three dollars you can get four pounds of potatoes. A proportion or ratio can be modeled as a fraction and in this case we can use 4 potatoes/$3 dollars. That is the ratio. We can figure out the cost of 1 pound of potatoes by setting up an equation where we set our ratio equal to 1/(x amount of dollars). 4/3 = 1/x We can cross multiply to get rid of the fractions within the equation. Cross multiplying looks like this: http://tinyurl.com/jjjj3dn So algebraically ours would look like 4 * x = 3 * 1 The result would be 4x =3 After we cross multiply we can solve the proportion like an equation. 4x =3 Divide by 4 x = 3/4 One pound of potatoes would be 3/4 of a dollar or $0.75 To check this solution we simply plug it back into the proportion 4/3 = 1/(3/4) It appears we have a complex fraction. To get rid of a complex fraction you just rewrite it as a division problem. That would look like this: 4/3 = (1/1) / (3/4) Remember when we divide fractions you take the reciprocal of second and multiply it by the first fraction. That would look like this: 4/3 = (1/1)*(4/3) 4/3 = 4/3 The solution is correct. The price for one pound of potatoes is $0.75. Part B is really easy. If the x coordinate is the pounds of potatoes and y coordinate is the price, what does this mean?

OpenStudy (triciaal):

Do you still need more help?

OpenStudy (jenessa_2413):

yes please i dont understand

OpenStudy (carrionbeast):

I tried my best to explain but I guess I wasn't good enough. Here try khan academy. https://www.khanacademy.org/math/algebra-basics/core-algebra-linear-equations-inequalities/ratios-core-algebra/v/writing-proportions

OpenStudy (jenessa_2413):

ok thank you

OpenStudy (triciaal):